# Non-Linear Scoring Model for Translation Quality Evaluation

### v2

@@ 7.5 Improvement of a Unified Calibrated Quality Scale
With a non-linear calibrated scoring model, it becomes feasible to normalize all evaluations to a unified, client-specific quality scale. Since the logarithmic tolerance accounts for sample-size effects, and calibration maps results to a common 0–100 (PT/MSV) scale, all evaluations land on the same company-wide quality scale—so scores from jobs of different lengths, language pairs and content types are directly comparable.
For example, replicating the English–German marketing case from Table 3 in [GladkoffEtAl2024], a 5,000-word translation with 23 minor errors passed the linear model, but was rejected by human reviewers. The logarithmic model set a threshold of 16 errors—correctly matching the expert decision.
@@ 7.6 Applications Beyond Human Translation
While the immediate application of this model is human translation quality evaluation, the underlying principles are equally relevant to assessing AI-generated content. Large language model (LLM) outputs—such as summaries, answers, or document translations—are increasingly being evaluated using scoring methods similar to those applied for human translation quality evaluation.
@@ 7.8 Summary
This model equips both translation professionals and tool developers with a framework that respects the cognitive and perceptual realities of language evaluation, enhancing decision-making at every level of the translation quality pipeline.
The proposed non-linear model captures how translation quality is actually perceived by humans-especially when evaluation sample sizes vary. It is simple to implement, easy to calibrate, and theoretically grounded. By adopting non-linear scoring, evaluation systems can better align numeric scores with human judgment, increasing the reliability and fairness of translation quality evaluations, as well as increasing the reliability and fairness of quality evaluations of translations created by humans or AI-generated texts.
@@ 8 Limitations and Future Work
The following limitations and research directions qualify the benefits summarized in §7.8 and the conclusion.
While the findings and proposed non-linear quality evaluation model presented in this paper are supported by both empirical observation and theoretical grounding, there are important limitations that must be acknowledged. Addressing these will be essential for refining the model and validating its broader applicability.
@@ 8.1 Limited Survey Group Size
We conducted structured interviews with quality managers at three very large enterprise clients, and the results are corroborated by their internal practices. Despite the scale and maturity of these organizations’ multilingual translation programs, the interview sample is still small. Additional evidence was drawn from our own practice, where we exclusively use a non-linear calibrated scoring model due to its universality and accuracy. However, the tolerance data we used was gathered through expert elicitation and professional experience, rather than under blind experimental conditions—introducing subjectivity and potential sources of bias, including response bias.
@@ 8.2 Content-Type Specific Effects
Different content types (e.g., marketing, legal, technical, UI strings) may have different quality expectations and error tolerance. Tolerance growth is assumed to follow the same logarithmic form $E(x)=a\,\ln(1+bx)$; the two coefficientsaa (overall scale) and $b$ (curvature)—may differ by client or use case.
We found that a single logarithmic formula performs well across a wide range of content types, suggesting that the logarithmic rule is quite universal—even with fixed parameters. However, this observation requires further validation through broader field data. A valuable extension of this work would be to investigate whether different genres (content types) exhibit distinct curve shapes—for example, steeper or flatter logarithmic profiles—or whether certain domains deviate significantly from the general trend. Such an analysis would help determine whether the proposed model serves as a general framework, as our experience suggests, or represents just one of several viable scoring paradigms.
We hypothesize a soft saturation at very large evaluation wordcount due to document‑level coherence expectations; future work can test whether $E(x)$ bends further (e.g., log‑log or logistic) at book/manual length.
@@ 8.3 Severity Weighting and Mixed Error Types
Severity scaling (i.e., fixed severity weights such as minor/major/critical multipliers) is incorporated by assigning a major or critical error a fixed multiple of the minor penalty points. This is separate from any error-type weights (e.g., terminology vs. grammar) a metric may also define. While a logarithmic curve fits minor-error tolerance well—minor errors being the most frequent and thus the most statistically stable—it may be less suitable for major errors, which occur far less often. For these rarer severities, an integer-valued sequence may be preferable, as indicated by Client 3’s data in Fig. 3. That sequence follows a loosely logarithmic trend but would require explicit rounding rules.
@@ 8.5 Extension to AI Evaluation Benchmarks
This paper also hypothesizes that the same principles of non-linear error perception apply to AI-generated text evaluation, including summarization, dialogue, and document translation by large language models. However, this claim has yet to be formally validated. AI benchmarks which rely on reference-based scoring or raw error density metrics can be improved by introducing calibrated non-linear models which could improve their alignment with human preferences.
@@ 9 Conclusion
This paper presents empirical data from experienced quality managers at large translation buyers, all of whom consistently demonstrated that their internal sense of acceptable error count rises non-linearly with sample size. This effect was evident across organizations, content types, and error categories and severities. This paper further supports these observations with theoretical underpinnings from psychophysics (via the Weber–Fechner law) and cognitive science (via Cognitive Load Theory), both of which explain why perceived quality and impact of errors do not scale linearly with text length.
@@
To address this disconnect between the existing scoring systems and human judgment, this paper proposes a calibrated logarithmic scoring model that reflects the genuine growth of error tolerance. This model is mathematically simple, empirically grounded, and easily adaptable to specific client requirements. It provides a flexible, yet rigorous alternative to traditional linear error density calculations.
Because it retains a simple two-parameter form—and requires only a one-dimensional solve for calibration—the model is straightforward to integrate into CAT tools, dashboards, and analytics, while providing a unified way to reflect how perceived impact scales with error accumulation.
@@
The implications of adopting this model are wide-ranging. Quality managers can gain a more accurate tool for defining passing thresholds in jobs of variable lengths. CAT tools and LQA systems and tools can offer context-aware scoring and fairer quality metrics. Evaluators benefit from consistency that reflects actual perception, and the same principles can be extended to evaluating AI-generated language output.
While further research is needed to validate and refine this model, particularly across languages, domains, content types, the fundamental insight seems to be clear: human perception of translation quality is inherently non-linear. By embracing this fact, the translation industry can take a critical step toward a more accurate, fair, and human-perception-aligned quality evaluation.
@@
Finally, this paper reiterates the warning from [GladkoffEtAl2024]: below the sample size of approximately 250 words, no deterministic tolerance curve is statistically reliable. In such micro-range cases, binomial risk-based Statistical Quality Control (SQC) remains the only viable method. [^7] Even with SQC, micro-range evaluations cannot provide statistically reliable estimates of the underlying population quality, because the confidence intervals around error density are too wide to be meaningful at such small evaluation sample sizes. SQC explicitly departs from the idea of estimating population quality and instead evaluates the producer’s risk (the risk of wrongly rejecting acceptable output) and the consumer’s risk (the risk of wrongly accepting defective output). In this way, SQC reframes quality control for very small samples: It does not reduce statistical uncertainty, but it makes risks explicit and quantifiable for decision-making.
This study consolidates the multi-range proposal of [GladkoffEtAl2024] by demonstrating that the same non-linear curve can be independently derived from psychophysical theory and empirical data, thus closing the theoretical gap identified in that earlier work.
Acknowledgments
@@
The authors thank the Quality Managers of the participating organizations (referred to in this document as Clients C1–C3) for their generous contributions of time, insight, and real-world data. Their candid responses and openness to critical reflection were essential to this research. Gratitude is also expressed to the broader translation quality evaluation community for ongoing discussions that continue to refine our shared understanding of what quality means in multilingual communication.
Appendix A. Calibration from Two Points (Numerical)
@@
Appendix B. Least-Squares Calibration from Multiple Points
@@
    If a feasible two‑point calibration is available ( Appendix A. Calibration from Two Points (Numerical)), use its $b^{\star}$ as the starting value; otherwise set $b_{0}=1/(10\,\max_{i}x_{i})$.
@@
Appendix C. Excel Goal Seek method for Two-Point Calibration
@@
Appendix D. Minimal Python code for Two-Point Calibration
@@
Appendix E. Least-Squares Fit (multiple points)
@@
Example 1 (Client 1, pages vs. minor errors; Fig. 3/Table 1).
@@
Appendix F. Wolfram Notebook for Least Squares Fit
@@
Illustration and Resulting Formula

### v3

@@ 7.5 Improvement of a Unified Calibrated Quality Scale
With a non-linear calibrated scoring model, it becomes feasible to normalize all evaluations to a unified, user-specific quality scale. Since the logarithmic tolerance accounts for sample-size effects, and calibration maps results to a common 0–100 (PT/MSV) scale, all evaluations land on the same company-wide quality scale—so scores from jobs of different lengths, language pairs and content types are directly comparable. For example, replicating the English–German marketing case from Table 3 in [GladkoffEtAl2024], a 5,000-word translation with 23 minor errors passed the linear model, but was rejected by human reviewers. The logarithmic model set a threshold of 16 errors—correctly matching the expert decision.
@@ 7.6 Applications Beyond Human Translation
While the immediate application of this model is human translation quality evaluation, the underlying principles are equally relevant to assessing AI-generated content. Large language model (LLM) outputs—such as summaries, answers, or document translations—are increasingly being evaluated using scoring methods similar to those applied for human translation quality evaluation [han2024neural].
@@ 7.8 Summary
This model equips both translation professionals and tool developers with a framework that respects the cognitive and perceptual realities of language evaluation, enhancing decision-making at every level of the translation quality pipeline. The proposed non-linear model captures how translation quality is actually perceived by humans-especially when evaluation sample sizes vary. It is simple to implement, easy to calibrate, and theoretically grounded. By adopting non-linear scoring, evaluation systems can better align numeric scores with human judgment, increasing the reliability and fairness of translation quality evaluations, as well as increasing the reliability and fairness of quality evaluations of translations created by humans or AI-generated texts.
@@ 8 Limitations and Future Work
The following limitations and research directions qualify the benefits summarized in Section 7.8 and the conclusion. While the findings and proposed non-linear quality evaluation model presented in this paper are supported by both empirical observation and theoretical grounding, there are important limitations that must be acknowledged. Addressing these will be essential for refining the model and validating its broader applicability.
@@ 8.1 Limited Survey Group Size
We conducted structured interviews with quality managers at three very large enterprise institutes, and the results are corroborated by their internal practices. Despite the scale and maturity of these organizations’ multilingual translation programs, the interview sample is still small. Additional evidence was drawn from our own practice, where we exclusively use a non-linear calibrated scoring model due to its universality and accuracy. However, the tolerance data we used was gathered through expert elicitation and professional experience, rather than under blind experimental conditions—introducing subjectivity and potential sources of bias, including response bias.
@@ 8.2 Content-Type Specific Effects
Different content types such as marketing, legal, technical, and UI strings, may have different quality expectations and error tolerance. Tolerance growth is assumed to follow the same logarithmic form $E(x)=a\,\ln(1+bx)$; the two coefficientsaa (overall scale) and $b$ (curvature)—may differ by user or use case.
We found that a single logarithmic formula performs well across a wide range of content types, suggesting that the logarithmic rule is quite universal—even with fixed parameters. However, this observation requires further validation through broader field data. A valuable extension of this work would be to investigate whether different genres (content types) exhibit distinct curve shapes—for example, steeper or flatter logarithmic profiles—or whether certain domains deviate significantly from the general trend. Such an analysis would help determine whether the proposed model serves as a general framework, as our experience suggests, or represents just one of several viable scoring paradigms. We hypothesize a soft saturation at very large evaluation wordcount due to document‑level coherence expectations; future work can test whether $E(x)$ bends further (e.g., log‑log or logistic) at book/manual length.
@@ 8.3 Severity Weighting and Mixed Error Types
Severity scaling, i.e., fixed severity weights such as minor/major/critical multipliers, is incorporated by assigning a major or critical error a fixed multiple of the minor penalty points. This is separate from any error-type weights (e.g., terminology vs. grammar) a metric may also define. While a logarithmic curve fits minor-error tolerance well—minor errors being the most frequent and thus the most statistically stable—it may be less suitable for major errors, which occur far less often. For these rarer severities, an integer-valued sequence may be preferable, as indicated by Institution 3’s data in Fig. 3. That sequence follows a loosely logarithmic trend but would require explicit rounding rules.
@@ 8.5 Extension to AI Evaluation Benchmarks
This paper also hypothesizes that the same principles of non-linear error perception apply to AI-generated text evaluation, including summarization, dialogue, and document translation by large language models. However, this claim has yet to be formally validated. AI benchmarks such as WMT [^10] https://machinetranslate.org/wmt shared task series [kocmi-etal-2025-findings] which rely on reference-based scoring or raw error density metrics can be improved by introducing calibrated non-linear models which could improve their alignment with human preferences.
@@ 9 Conclusion
This paper presents empirical data from experienced quality managers at large institutional translation service users, all of whom consistently demonstrated that their internal sense of acceptable error count rises non-linearly with sample size. This effect was evident across organizations, content types, and error categories and severities. This paper further supports these observations with theoretical underpinnings from psychophysics (via the Weber–Fechner law) and cognitive science (via Cognitive Load Theory), both of which explain why perceived quality and impact of errors do not scale linearly with text length.
@@
To address this disconnect between the existing scoring systems and human judgment, this paper proposes a calibrated logarithmic scoring model that reflects the genuine growth of error tolerance. This model is mathematically simple, empirically grounded, and easily adaptable to specific user requirements. It provides a flexible, yet rigorous alternative to traditional linear error density calculations. Because it retains a simple two-parameter form—and requires only a one-dimensional solve for calibration—the model is straightforward to integrate into CAT tools, dashboards, and analytics, while providing a unified way to reflect how perceived impact scales with error accumulation.
@@
The implications of adopting this model are wide-ranging. Quality managers can gain a more accurate tool for defining passing thresholds in jobs of variable lengths. CAT tools and LQA systems and tools can offer context-aware scoring and fairer quality metrics. Evaluators benefit from consistency that reflects actual perception, and the same principles can be extended to evaluating AI-generated language output. While further research is needed to validate and refine this model, particularly across languages, domains, content types, the fundamental insight seems to be clear: human perception of translation quality is inherently non-linear. By embracing this fact, the translation industry can take a critical step toward a more accurate, fair, and human-perception-aligned quality evaluation.
@@
Finally, this paper reiterates the warning from [GladkoffEtAl2024]: below the sample size of approximately 250 words, no deterministic tolerance curve is statistically reliable. In such micro-range cases, binomial risk-based Statistical Quality Control (SQC) remains the only viable method. [^11] Even with SQC, micro-range evaluations cannot provide statistically reliable estimates of the underlying population quality, because the confidence intervals around error density are too wide to be meaningful at such small evaluation sample sizes. SQC explicitly departs from the idea of estimating population quality and instead evaluates the producer’s risk (the risk of wrongly rejecting acceptable output) and the consumer’s risk (the risk of wrongly accepting defective output). In this way, SQC reframes quality control for very small samples: It does not reduce statistical uncertainty, but it makes risks explicit and quantifiable for decision-making. This study consolidates the multi-range theory findings by demonstrating that the same non-linear curve can be independently derived from psychophysical theory and empirical data, thus closing the theoretical gap identified in that earlier work.
10 acknowledgments
@@
The authors thank the Quality Managers of the participating organizations (referred to in this document as users C1–C3) for their generous contributions of time, insight, and real-world data. Their candid responses and openness to critical reflection were essential to this research. Gratitude is also expressed to the broader translation quality evaluation community for ongoing discussions that continue to refine our shared understanding of what quality means in multilingual communication.
Appendix A Institution User Quotations on Evaluation
A recent quote from a large institutional user illustrates this point:
“Once we started using our current methodology in 2020, we still asked the evaluators to indicate the cases where their actual feeling was different from what the score gave them. We very quickly realized that the main issue was that with very short samples the scoring was overly harsh and with very long samples it was too lenient. The reason for this is that when we evaluate holistically, the perception is not congruent with our scoring formula. For example, we might feel that if a translation sample is about one page, a single major mistranslation error is already enough to judge it as failing; in a seven-page sample, however, seven such errors are far more than we would be willing to tolerate. Instead, we would prefer to fail the sample already at three or four major errors. This poses a problem for the linear scoring model which simply prorates the number of errors per page to a total number of pages in the sample.”
Another industrial user confirmed these findings and chose to avoid the issue by only evaluating samples of exactly 1000 words. However, in real-world operations, such precisely consistent evaluation sample sizes are rare. Evaluation samples often vary significantly in length due to the nature of the content, deadlines, or operational constraints.
Appendix B Linear vs Logarithmic Fits
In other words, the visual impression is also statistically decisive. Below is the quantitative comparison of linear vs. logarithmic fits on the elicited points. To decide whether a straight line or a logarithmic curve better represents the expert data, we fit both forms to the same points and then measure how close each comes. This is what the values in Table 3 show: SSE and RMSE are the remaining gaps between curve and data (smaller is better), $R^{2}$ is the share of variation explained (closer to 1 is better), and AIC/BIC compare models while penalizing complexity (lower is better). Even after each curve is fitted as well as possible to the data, the log model has dramatically smaller errors, much higher $R^{2}$, and far lower AIC/BIC. This means that the shape of the logarithmic curve matches the human judgments far more closely than the linear line ever can. In other words, it is not just that we chose to use a logarithm; the data themselves show that the logarithmic shape is the right one.
Model $a$ $b$ $c$ SSE RMSE $R^{2}$ AIC BIC
Linear (through origin) cxcx $0.541$ $26.755$ $1.955$ $0.044$ $11.386$ $11.331$
Table 3: Goodness-of-fit statistics for the user 1 minor-error dataset ($x$ in pages; $E$ in minor points). Lower SSE/RMSE/AIC/BIC and higher $R^{2}$ indicate better fit. Data points: (2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8)(2,2),(3,3),(4,4),(5,5),(7,6),(10,7),(20,8).
Appendix C Calibration from Two Points (Numerical)
@@
Appendix D Least-Squares Calibration from Multiple Points
@@
    If a feasible two‑point calibration is available ( Calibration from Two Points (Numerical)), use its $b^{\star}$ as the starting value; otherwise set $b_{0}=1/(10\,\max_{i}x_{i})$.
@@
Appendix E Excel Goal Seek method for Two-Point Calibration
@@
Appendix F Minimal Python code for Two-Point Calibration
@@
Appendix G Least-Squares Fit (multiple points)
@@
Example 1 (user 1, pages vs. minor errors; Fig. 3/Table 1).
@@
Appendix H Wolfram Notebook for Least Squares Fit
@@ H.1 Instructions
H.2 Illustration and Resulting Formula